# Not feeling the buzz: Correction study of mispricing and inefficiency in online sportsbooks

### v3

@@ 4.3 Bookmaker mispricing
Finally, we tested bookmaker mispricing by replicating the original mispricing test (Ramirez et al., 2023, Table 1, Column IV) on the preferred model (PM) over the period Feb. 2020 – Aug. 2023. Shown in Table 4, we find that new estimated coefficients for WikiBuzz and RankDist are no longer significant over this period, with respective p-values of $0.650$ and $0.531$.
@@
Odds-implied probability, $\hat{\beta_{1}}$ $-0.043$ $0.188$
@@
WTA rank distance to opponent (RD), $\hat{\beta_{2}}$ $0.018$ $0.531$
@@
Wiki relative buzz factor, $\hat{\beta_{3}}$ $0.002$ $0.650$
@@ 6 Conclusions
Our work exemplifies the sensitivity of betting models to data quality, and expands upon the findings of Wunderlich and Memmert (2020) to demonstrate that even a single data error can significantly influence ROI when adopting a Kelly staking strategy. Therefore, simulated betting returns should be interpreted cautiously, and data quality should always be checked, even when using popular open soure datasets such as tennis-data.co.uk.
@@ Appendix B Best Odds Strategy Results from Mispricing Test
Betting results using best odds are almost exactly replicated, apart from a minor discrepancy of $N=5,188$ odds (Ramirez et al., 2023, Table 4, Column (II)) as opposed to the $N=5,189$ odds used in our replication. We traced the discrepancy in $N$ odds of $1$ to a calculation error by RRS regarding the match between Anastasia Potapova and Kristina Kucova on January 15, 2019.
The best odds for Kucova are unavailable, but they are available for Potapova. When calculating the $N$ odds used, the logic used by Ramirez et al. (2023) doesn’t count either bet. However, to obtain the results shown in Table 6 the Potapova bet must be included, as it results in a difference in ROI of $0.02$. Therefore, the true value of $N=5189$.
@@ 4. Best odds:
The final dataset contains 35,274 rows (N odds of player-matches), covering 17,751 tennis matches and 584 players. Match dates range from 02 July 2016 to 31 August 2023. Cleaned data is available online. [^17] Cleaned data: https://github.com/Faxulous/notFeelingTheBuzz.

### v4

@@ 4.3 Bookmaker mispricing
Finally, we tested bookmaker mispricing by replicating the original mispricing test (Ramirez et al., 2023, Table 1, Column IV) on the preferred model (PM) over the period Feb. 2020 – Aug. 2023. Shown in Table 4, we find that new estimated coefficients for WikiBuzz and RankDist are no longer significant over this period, with respective p-values of $0.837$ and $0.471$.
@@
Odds-implied probability, $\hat{\beta_{1}}$ $0.039$ $0.148$
@@
WTA rank distance to opponent (RD), $\hat{\beta_{2}}$ $0.021$ $0.471$
@@
Wiki relative buzz factor, $\hat{\beta_{3}}$ $0.001$ $0.837$
@@ 6 Conclusions
Our work exemplifies the sensitivity of betting models to data quality, and expands upon the findings of Wunderlich and Memmert (2020) to demonstrate that even a single data error can significantly influence ROI when adopting a Kelly staking strategy. Therefore, simulated betting returns should be interpreted cautiously, and data quality should always be checked, even when using popular open source datasets such as tennis-data.co.uk.
@@ Appendix B Best Odds Strategy Results from Mispricing Test
Betting results using best odds are almost exactly replicated, apart from a minor discrepancy of $N=5,188$ odds (Ramirez et al., 2023, Table 4, Column (II)) as opposed to the $N=5,189$ odds used in our replication. We traced the discrepancy in $N$ odds of $1$ and mean overround of $0.01$ to a calculation by RRS regarding the match between Anastasia Potapova and Kristina Kucova on July 15, 2019.
The best odds for Kucova are unavailable, but they are available for Potapova. When calculating the $N$ odds used, the logic used by Ramirez et al. (2023) doesn’t count either bet. However, to obtain the results shown in Table 6 the Potapova bet must be included, as it results in a difference in ROI of $0.02$. Therefore, the true value of $N=5189$ and the mean overround is $-0.24\%$.
@@ 4. Best odds:
The final dataset contains 35,274 rows (N odds of player-matches), covering 17,751 tennis matches and 584 players. Match dates range from 03 July 2015 to 31 August 2023. Cleaned data is available online. [^17] Cleaned data: https://github.com/Faxulous/notFeelingTheBuzz.